OpenStudy (anonymous):

The measure of an angle is one third the measure of its supplement. What is the measure of the angle?

OpenStudy (bibby):

If we have an angle, let's say A, how would you express its supplement?

OpenStudy (anonymous):

Would you just divide the angle by 1/3 @bibby

OpenStudy (pawanyadav):

Hint:sum of supplementry Angles is 180°

OpenStudy (bibby):

if we call the supplementary angle B, we have A+B=180

OpenStudy (anonymous):

But how do you figure out what A is? @bibby

OpenStudy (pawanyadav):

Now you can do B=1/3(A)

OpenStudy (pawanyadav):

Then plug this value of B in A+B=180° And you get equation in single variable AA so you can find it easily

OpenStudy (pawanyadav):

In variable A

OpenStudy (pawanyadav):

@kenna_098 is this helping you

OpenStudy (anonymous):

I'm still really confused @Pawanyadav

OpenStudy (pawanyadav):

A+B=180 and B=A/3 So A+A/3=180

OpenStudy (pawanyadav):

Can you find A now

OpenStudy (mrnood):

 sum of these 2 angles = 180 (because they are 'supplementary')